OpenStudy (anonymous):

What is the extra step that must be taken when dealing with rational exponents that you do not have to worry about with integer exponents?

OpenStudy (anonymous):

@ilfy214 can you help

OpenStudy (zzr0ck3r):

(-2)^2 is fine \[(-2)^{\frac{1}{2}}=\sqrt{-2}\] is not

OpenStudy (anonymous):

so it would be the neg number

OpenStudy (anonymous):

changing it to a pos.?

OpenStudy (anonymous):

@zzr0ck3r

OpenStudy (anonymous):

no you don't just change it to positive

OpenStudy (zzr0ck3r):

\[a^{\frac{b}{c}}\] c cant be even when a <0

OpenStudy (zzr0ck3r):

unless b is even

OpenStudy (zzr0ck3r):

you cant take the an even root of a negative number

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so when the denominator is a pos u gotta change it to a neg. ???

OpenStudy (anonymous):

you wouldnt have to worry about changing it to a pos

OpenStudy (zzr0ck3r):

I said nothing about a denominator being positive \[a^{\frac{b}{c}}\\ if\space a<0 \space and \frac{b}{c}\space is\space in\space lowerst\space terms\] then c cant be EVEN

OpenStudy (anonymous):

ok